OpenStudy (rational):

Show that the product of any 3 consecutive positive integers is divisible by 6 using Fermat's infinite descent

OpenStudy (ikram002p):

What method is this :(

OpenStudy (anonymous):

It seems to me like there is no useful way for infinite descent method in integer domain :( The statement would look like this: "There is no smallest n for which the product is divisible." But it always drops down to standard divisibility proof - p|2 and p|3 -> p|6.

OpenStudy (anonymous):

I do not know what method this is, but logically, if you have 3 consecutive integers, at least 1 of these numbers will be divisible by 2 and another will be divisible by 3, hence why they are also divisible by 6.

OpenStudy (anonymous):

I mean if n is not divisible by 2 then n+1 is divisible by 2 and if n is also not divisible by 3 then n+2 is divisible by 3. If n is divisible by 3 then n+1 is divisible by 2.

OpenStudy (rational):

@Frouzen Yes I think any kind of proof with infinite descent must reach below step "6 does not divide 1*2*3" which is a contradiction, thus ending the proof.

OpenStudy (rational):

@zimmah It is easy to see that the product of any two consecutive integers is divisible by 2 because one of them is always even. But how do you prove divisibility by 3 ?

OpenStudy (anonymous):

@rational Well, think about it, take any number (n) at all, and see if it divides by 3. If it doesn't then either n+1 or n+2 MUST divide by 3.

OpenStudy (anonymous):

with number i mean integer.

OpenStudy (anonymous):

odd and even numbers are usually described as 2k and 2k+1, k integers, where 2k+2 can be written as 2(k+1) the same for division by 3: 3k 3k+1 3k+2 3k+3 = 3(k+1) therefore the proof needs to be made for the three possibilities

OpenStudy (rational):

Nice! :) Frouzen is that for infinite descent ?

OpenStudy (anonymous):

3k: n(n+1)(n+2) = 3k(3k+1)(3k+2) = 3 [k(3k+1)(3k+2) 3k+1: (3k+1)(3k+2)(3k+3) = (3k+1)(3k+2)·3(k+1) = 3 [(3k+1)(3k+2)(k+1)] 3k+2 simillary. That is the standard proof of many divisibility problems. But I cannot see any logical way how to fit the infinite descent in...

OpenStudy (rational):

Ahh yes division algorithm always works nicely for these divisibility problems when the modulus is small

OpenStudy (rational):

I'm thinking something like this - With the idea of arriving at a contradiction, let us assume that there exists a product of consecutive integers \(n(n+1)(n+2)\) that is not divisible by \(6\) : \[6\nmid n(n+1)(n+2)\] From this, we need to construct another smaller product of consecutive positive integers thats not divisible by \(6\). Since \(2\mid n(n+1)\) gives \(6\mid 3n(n+1)\) we have : \[\begin{align}&6\nmid[n(n+1)(n+2)-3n(n+1)]\\ &6\nmid n(n+1)(n+2-3)\\ &6\nmid n(n+1)(n-1)\\ &6\nmid (n-1)n(n+1)\\ \end{align}\] which is another smaller product of consecutive integers not divisible by \(6\). repeating the whole argument we see that \[\begin{align}&6\nmid (n-2)(n-1)n\\ &~~~~~~~\vdots\\ &~~~~~~~\vdots\\ &6\nmid 1\cdot 2\cdot 3\\ \end{align}\] This is a contradiction because \(6\mid 6\), and ends the proof.

OpenStudy (anonymous):

I would go for it this way. Actualy, I see no reason why not to go for (n-2)(n-1)n from the beginning as it is also a product of three consecutive integers, but this looks more... mathematically ... :D

OpenStudy (rational):

I thought of using something else for 3n(n+1) but couldn't find anything more simpler... thanks for checking, appreciate it very much :)

OpenStudy (ikram002p):

So its sort of combination btw well ordering principles and proof by contradiction ??

OpenStudy (ikram002p):

I should learn that :)

OpenStudy (rational):

Yes it is a contradiction proof using induction